One pipe can fill a tank in 4 hours. A second pipe also requires 4 hours but a third needs three hours. How long will it take to fill the tank if all three pipes are open?
Answer by richwmiller(17219)

1/4+1/4+1/3=1/x
1/2+1/3=1/x
3/6+2/6=1/x
5/6=1/x
6/5=x
1.2=x